OpenStudy (animalover02):

Anyone want to help with 5 math questions? ill medal & fan!!!!

OpenStudy (m.d.saxon):

hi

OpenStudy (m.d.saxon):

what are they

OpenStudy (animalover02):

one sec..ill post one

OpenStudy (animalover02):

nancy bought 360 tropical fish for a muesem display. she bought 8 times as many goldfish as angelfish. How many of each type of fish did she buy? Which system of equation models the problem if x represents the number of angelfish nancy bought and y represents the number of goldfish she bought? A. x-y=360 y=8x B. x+y=360 y=8x C. x+y=8 y=360x D. x+8y=360 y=8x

OpenStudy (m.d.saxon):

its either A or B but i cant determine which one